OpenStudy (anonymous):

solve for x x^2+6x=46

OpenStudy (anonymous):

put in standard form: x^2 + 6x - 46 = 0 now use the quadratic formula....

OpenStudy (anonymous):

-6+/-sqrtof220/6?

OpenStudy (phi):

1st, you divide by 2*a (a being the coefficient of the x^2 term) a = 1 here so divide by 2 not 6 2nd, you should simplify sqrt(220) factor 220 into its prime factors 2*110 2*11*10 2*2*5*11 there is a pair of 2's so you can rewrite sqrt(220) as sqrt(2*2)*sqrt(55)= 2sqrt(55) the answer is \[ \frac{-6±2\sqrt{55}}{2}= -3±\sqrt{55} \]